### v1

# A flexible quantum data bus

Abstract
We consider multi-path routing of entanglement in quantum networks, where a pre-prepared multipartite entangled 2D cluster state serves as a resource to perform different tasks on demand. We show how to achieve parallel connections between multiple, freely chosen groups of parties by performing appropriate local measurements among diagonal paths, which preserves the entanglement structure of the remaining state. We demonstrate how to route multiple Bell-states along parallel lines via crossings, turns and fade-in/-outs, analogously to a data bus. The results apply to networks at any scale.
preprint: APS/123-QED
Introduction.—
Connecting devices constitutes the foundation of modern society. All devices operate in some form of network, with the goal of either increasing computational power, to perform distributed tasks or to simply exchange messages of some kind. The structure and implementations of such networks often depend on the sizes of devices. For example, for embedded devices such a network corresponds to connected, hard-wired micro-controllers. In contrast computers connect through local-area-networks over WiFi. With the advent of quantum computers approaching, researchers believe the next evolutionary step in the computing industry lies in the connection of quantum computing devices, with the ultimate goal to unleash new applications impossible by classical means. Such applications include quantum key distribution [1], secret sharing [2, 3], sensor networks [4, 5] but also distributed quantum computing [6, 7, 8].
At the lowest level, quantum devices, in any quantum network [9, 10, 11, 12, 13], at any scale, are connected via physical channels. A Bell-state between nodes is equivalent to a channel due to the teleportation protocol [14], paving the way for sharing and manipulating entanglement [15] as an alternative to channels. In recent years, the vast majority of research on quantum networks has focused on the generation of Bell-states between two nodes in a network of Bell-states [16, 17, 18, 19, 20, 21, 22, 8, 13, 23, 24, 25]. This generation involves finding an optimal path [26] (in terms of some metrics) within the network and connecting the states along that path. Here, we introduce a flexible quantum data bus by establishing multiple Bell-states simultaneously in a network of any scale by utilizing pre-established 2D cluster states as resource, referred to as cluster states [27, 28] in the following, instead of a network of Bell-states as a resource for routing [29, 30].
In our setting, devices may connect at any node in the cluster state, which is generated as a flexible resource state to fulfill different tasks on demand, even prior to network requests. The manipulation of resource states takes place solely by single-qubit measurements and single-qubit unitaries. At the heart of our proposal is the zipper-scheme, a measurement pattern among diagonal lines to establish a Bell-state within the cluster state. We remark that the zipper-scheme is based on the X-protocol introduced in [31]. Other works also study staircase-shaped measurement patterns on 2D cluster states, even demonstrating optimality with regards to the number of measurements [32, 33]. As we demonstrate, the zipper-scheme preserves the entanglement structure of the remaining cluster state when generating a Bell-state. Keeping the grid structure intact and usable turns out to be pivotal to extract further Bell-states from the cluster state. We utilize this remarkable property to establish multiple Bell-states between arbitrary nodes in a cluster state, including crossings and turns of the measurement paths. We note, however, that there exist some prohibited areas around the turning and end points, where Pauli Z measurements are necessary to cut out the Bell-state from its direct neighborhood. However, this is still in stark contrast to standard methods [34, 28, 35], where the connection between nodes is established by cutting a hole among the entire path. This method not only wastes resources, but also hinders the generation of further Bell-states. Our scheme enables to simultaneously create Bell-states along adjacent measurements paths, effectively constructing a quantum data bus in analogy to classical busses found for instance in mainboards. We compose a quantum data bus architecture in a modular fashion by combining elementary building blocks such as crossings, turns and fade-in/-outs. Our results have a multitude of applications from which we discuss its applicability ranging from large scale networks such as entanglement-based quantum networks  [29, 30, 36, 37] to local networks relying on a central unit [38, 39]. Furthermore, we believe that also small and integrated networks of quantum devices, similar to devices running in embedded environments like cars, benefit from a dynamic way of creating Bell-states on demand, as this removes the necessity to multiplex multiple connections over a single transmission channel like in classical, hard-wired controller networks. Diagonal routing also naturally extend to the multiparty case, like for example establishing GHZ states between multiple nodes within the cluster.
Setting.—
We consider quantum networks at different scales with the goal of establishing entanglement between multiple pairs of nodes in parallel. Rather than generating Bell pairs directly, we consider an entanglement-based scenario where nodes share a pre-prepared multipartite entangled resource state which they manipulate by means of measurements to establish desired configurations in a flexible way on demand. Such an approach has the advantage that it reduces latency, and allows to optimize the entanglement topology [29, 40, 41]. On the contrary it requires a long time memory for storage. Here we consider a 2D cluster state as a resource, where each node holds a single qubit of the cluster state. Such 2D cluster states serve as universal resource in measurement-based quantum computation (MBQC) [34, 28, 42], where information is processed by means of single qubit measurements only and arbitrary target states can be generated in this way. Here, in contrast, our sole goal is to establish multiple Bell pairs, which implies that considerations such as information flow [43, 44] do not apply, and additional methods and techniques are available. In particular, we are interested in establishing (multiple) Bell pairs between arbitrary nodes, independent of their location within the cluster state.
Preliminaries.—
Graph states [45, 46] are pure, multipartite entangled quantum states described by a classical graph $G=(V,E)$, where the vertex set $V$ and edge set $E$ correspond to qubits and binary Ising-type interaction between those, respectively. Formally, a graph state |$\ket{G}$ is uniquely defined as the eigenvector with $+1$ eigenvalues of the set of stabilizers
$K_{a}=\sigma_{x}^{a}\sigma_{z}^{N_{a}}=\sigma_{x}^{a}\prod_{b\in N_{a}}\sigma_{z}^{b}$ (1)
for each vertex $a\in V$ with its corresponding neighbors $N_{a}\subseteq V$. The dynamic picture is an alternative mean to define graph states, where all qubits are initially prepared in the $\ket{+}$ state and according to the edge set controlled-phase gates are applied between two vertices.
In this work we frequently use local complementation (LC), a specific local Clifford unitary, to transform a graph state. This unitary inverts the sub-graph induced by the neighbors of a qubit. Local Pauli Z measurements translate to vertex and edge erasure of the measured qubit. A Pauli Y measurement corresponds to a LC followed by a Pauli Z measurement on the qubit to measure. Local Pauli X measurements are a combination of LC on a specific neighbor qubit, a Pauli Y measurement on the qubit itself and again LC on the specific neighbor, see Ref. [46].
Zipper-scheme.—
The zipper-scheme tackles the problem of connecting two nodes in a 2D cluster state via two diagonal lines, thereby generating a Bell-state. The scheme accomplishes this by performing Pauli X measurements along a staircase-shaped measurement path (as introduced in [31] as X-protocol), where two such paths are combined to connect any two points on the grid. For a direct diagonal connection, one of the Bell-state qubits is used as neighbor for LC, and Pauli Z measurements in the neighbourhood of the endpoints isolate the final Bell-state. The inset in Fig. 1 illustrates the result of the zipper-scheme after the orange qubits have been measured in the X basis along a diagonal, whereas the purple qubit is one of the Bell-state constituents and has been used as reference qubit for LC. We observe that the zipper-scheme restores the remaining cluster state structure, see the red edges in Fig. 1, up to holes at the endpoints. Note that some Pauli Z basis measurements are necessary to fully isolate the Bell-state though, see the yellow qubits in Fig. 1. We investigate the properties of the protocol in depth in the Appendix.
Pauli X measurement
Pauli Z measurement
Qubit measured in Pauli X basis
Figure 1: Zipper-scheme (inset): The orange qubits are measured in the Pauli X basis such that the purple qubit is routed along the diagonal measurement path while the obtained edges between the red qubits restore the underlying cluster state. The qubits in the neighborhood of the purple qubit need to be removed by Pauli Z measurements leading to holes. Simultaneous Bell-state routing (main figure): We extract three Bell-states, purple, blue and turquoise, by applying the zipper-scheme along diagonals of orange qubits. Merely on turning and end points holes appear due to the yellow Pauli Z measurements to disconnect the Bell-state qubits from the remaining cluster state. The zipper-scheme enables crossing of paths except in the vicinity of end and turning points, and note that the paths are adjusted such that the measurement sequence is taken into account (first purple, then turquoise and blue).
The zipper-scheme allows us to generate a Bell-state on a diagonal line within the cluster state. However, we observe, as explained in more details in the Appendix (note that routing a single Bell-state is a special case of routing parallel Bell-states), that it is also possible to use the very same scheme to route a Bell-state around corners (’L’-turn), upside down (’V’-turn), diagonal-to-straight or over a crossing with another Bell-state. These ingredients enable us to route Bell-states of arbitrary topology on the cluster state. In Fig. 1 we demonstrate an example in which three Bell-states among arbitrary nodes are generated. We emphasize that two points on a 2D lattice can be always connected by two (or more) diagonals as shown in the Fig. 1. Importantly, the zipper-scheme allows to establish crossings of Bell-states on the cluster state provided the crossings are sufficiently far away from the end or turning points. Note that we do not sketch the additional edges after applying the zipper-scheme to simplify readability, and furthermore, we have adjusted the measurement paths assuming that the purple path is measured first followed by the turquoise and blue one. The ability to support crossings has two advantages. First, it prevents potentially long measurement paths around obstacles in a cluster state, and second, it allows to have simultaneous request also in cases that require crossings due to resource constraints, impossible to achieve without them. Using a 2D cluster state of size $n\times n$ allows one to obtain $\mathcal{O}(n)$ parallel Bell pairs. This can be seen by observing that any two points in a 2D cluster are connected by diagonal paths of length at most $2n$, which also corresponds (up to few additional measurements around end- and turning points) to the number of qubits that need to be measured to establish a Bell pair. Since the underlying entanglement structure is still intact, in total $\mathcal{O}(n)$ Bell pairs can be generated generically. Notice that some configurations may not be possible, as small holes are cut in the 2D structure around end- and turning points.
Quantum data bus.—
The next step is to extend the results of the zipper-scheme and building blocks to route multiple Bell-states along parallel lines and to which we refer to as a quantum data bus. We note that a quantum data bus is conceptually closely related to data busses in conventional computers used for connecting individual microchips with each other. A quantum data bus uses the entanglement of a cluster state in an efficient way to achieve that goal. In Fig. 2 we illustrate the main idea.
Figure 2: Quantum data bus and its building blocks: The red lines represent parallel transport along diagonal lines, including crossings orthogonal to the diagonals. The black lines show a parallel transport, both vertical and horizontal. The blue and pink lines depict parallel ’L’- and ’V’-turns, whereas the green lines demonstrate a fade-in/-out of data lines. Finally, the yellow qubits depict qubits on which we need to perform Pauli Z measurements.
Similar when routing a single Bell-state we identify multiple building blocks for routing Bell-states in parallel. Five of the parallel building blocks naturally extend the building blocks for a single Bell-state to multiple parallel Bell-states (diagonal, crossing, ’L’-turn, ’V’-turn, diagonal-to-straight). For instance, from a cluster state of size $n\times n$, it is possible to generate $\mathcal{O}(n)$ parallel Bell-states as part of a crossing. The required number of qubits to turn $n$ data lines in an ’L’-turn around the corner is $n\times n$. For ’V’-turns and the building block that turns diagonal Bell-states into a parallel Bell-states, the number of data lines $n$ of the building block determines the minimum, necessary length of the zipper-scheme, corresponding to $2n$ for the first data line. Therefore, it requires in total $n\times n$ measurements.
A new building block corresponds to parallel transport among a straight line. We achieve this kind of transport by applying the zipper-scheme repeatedly. Specifically, first one applies the zipper-scheme from the first Bell-state line towards the last line. This effectively routes the first Bell-state diagonally to the last line. Next, one applies the same scheme to the second Bell-state, moving it to the second last line. Repeatedly applying this pattern results in a total permutation of all lines. Performing the same scheme another time inverts the permutation, resulting in straight lines again. However, it should be noted that additional Pauli Z measurements are necessary to isolate the established data bus. A single qubit line is kept between the two permutation measurements, because the zipper-scheme requires to have a cluster state. In total it requires a block of length $2n$ qubits to transport $n$ data lines in parallel, which induces a fixed, minimal block size. Notice, though that there is no distance between the data lines within the data bus, in stark contrast to standard data transport schemes in MBQC [34, 28, 35] which require to isolate a path, resulting in a distance of one between two data lines, i.e. half the capacity as in our scheme.
Another new building block allows to fade-in or -out data to the data bus. It uses Pauli Z and X measurements to separate and merge the data lines, respectively. For adding a single data line we require one Pauli Z and one Pauli X measurement to cut and merge the line into the data bus.
We discuss all building blocks and its measurements in the Appendix in more detail. We also point out that diagonal routing straightforwardly extends to GHZ states, as described in the Appendix. Essentially, one can connect several qubits to some central one via different (diagonal) paths, thereby establishing a GHZ state.
Applications.—
We have laid the theoretical foundation for routing Bell-states on a 2D cluster state. In the following we discuss three potential use-cases of the parallel zipper-scheme, highlighting its applicability and importance.
The first application corresponds to a long-distance quantum network in which the nodes connect via a 2D cluster state. This setting was e.g. studied in [47, 48], in which the authors presented ways how to create a long-distance cluster state by purifying and merging small buildings blocks. Once the cluster state is established, the goal of the network is to generate Bell-states between any of the nodes in the cluster state on demand, maybe even in parallel. Using the results of this work we find that it is possible to connect any, sufficiently far, sets of communication partners via Bell-states by using the parallel zipper-scheme.
But also centralized communication scenarios with a powerful central unit as studied in [38, 39] benefit from our results. For that purpose, suppose that the nodes of a small scale network, similar to a LAN, connect to a central unit via (potentially multiple) Bell-states. In case the central unit uses internally a local cluster state, and the parties connect via a Bell-state to the border of the cluster state, the central unit is able to establish arbitrary, even parallel, connections amongst the nodes by using our quantum data bus proposal. In other words, the central unit uses the internal cluster to permute connections amongst the (outside) nodes. This allows for a flexible way to operate the central unit, as this unit is capable of connecting any of the connected nodes as necessary and on demand.
The last application we highlight corresponds to integrated quantum devices. In this vision, small quantum devices like sensors connect within the integrated quantum device via a Bell-state to the border of a 2D cluster state, similar in spirit to the previous application. This scenario is inspired by classical computing architectures in which microcontrollers connect through busses in a fixed manner. Our results imply that for quantum computing architectures such a fixed wiring is not necessary, but can be established on-demand in terms of quantum data busses by consuming a cluster state. This introduces flexibility, but also extend-ability to integrated quantum device design.
Conclusion.—
In this letter we have considered pre-established 2D cluster states as resource for quantum networks. We identified the main ingredient to route Bell-states in a cluster state, namely the zipper-scheme. The scheme preserves the cluster state entanglement structure, vital for establishing crossings of Bell-states on a cluster. It further turns out that the zipper-scheme lies at the heart of many building blocks to route Bell-states in a cluster state, like for example ’V’-turns or ’L’-turns. We also demonstrated how to further optimize these building blocks by showing how to utilize them for parallel transport of Bell-states. These building blocks play a key role for routing multiple Bell-states in a long-distance network, but also for small scale networks and the design of integrated quantum processing devices. We also discussed an extension of the diagonal routing to multiparty states like for example GHZ states, and we leave the extension of the other building blocks for future work. As an outlook, it remains an open question how noise in the resource state and the apparatus influences the fidelity of the routed Bell-states. First results for a single target Bell pair [32] indicate robustness against imperfections.
Acknowledgments
This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/P36010. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
Appendix A Graph states
Graph states [45, 46, 27] are a specific class of pure, multipartite entangled quantum states whose correlations correspond to a classical graph $G=(V,E)$. The vertex set $V$ and the edge set $E$ correspond to qubits and binary Ising-type interaction between qubits, respectively. Fig. 3 depicts an example graph state in the center. The figure demonstrates how the graph state transforms under certain local operations, like for example local complementations or Pauli basis measurements. Formally, a graph state |$\ket{G}$ with respect to the graph $G$ is constructed from the set of qubits in the $\ket{+}$ state through the interaction
|$\ket{G}=\prod_{\{a,b\}\in E}C\mathrm{Z}_{ab}\ket{+}^{\otimes|V|},$ (2)
where a controlled-Z gate defined as $diag(1,1,1,-1)$ is applied, if the edge between the qubits $a$ and $b$ exits in the graph $G$. For example the central graph in Fig. 3 comprises four qubits, labeled from one to four. One generates the corresponding graph state by applying controlled-Z gates between the qubits one and two, one and four, one and three, two and three, and two and four, respectively.
Figure 3: In the center is a graph state depicted, whereas vertices correspond to qubits, which are initialized in the $\ket{+}$ state and edges represent maximally entanglement between two qubits. In the corners the resulting graph states are shown, which are caused by local single-qubit Pauli measurements or local complementation as indicated by the labeled arrows.
Alternatively, a graph state |$\ket{G}$ is defined as the common, unique eigenstate with eigenvalue $+1$ of the set of stabilizers
$K_{a}=\sigma_{x}^{a}\sigma_{z}^{N_{a}}=\sigma_{x}^{a}\prod_{b\in N_{a}}\sigma_{z}^{b},$ (3)
for each vertex $a$ with its corresponding neighbors $N_{a}$. Consequently, we describe the central graph in Fig. 3 uniquely via the stabilizers $K_{1}=\sigma_{x}^{1}\sigma_{z}^{2}\sigma_{z}^{3}\sigma_{z}^{4}$ and $K_{2}=\sigma_{x}^{2}\sigma_{z}^{1}\sigma_{z}^{3}\sigma_{z}^{4}$, $K_{3}=\sigma_{x}^{3}\sigma_{z}^{1}\sigma_{z}^{3}$ and $K_{4}=\sigma_{x}^{4}\sigma_{z}^{1}\sigma_{z}^{2}$.
A.1 Local Clifford unitaries and Pauli measurements
Local transformations of graph states play a crucial role in distributed settings. Many of them correspond to graphical manipulations rules on the classical graph $G$, including local, single-qubit Clifford operations and Pauli measurements. For example, local complementation corresponds to applying the unitary operator $\sqrt{K_{a}}$ to the graph state |$\ket{G}$. Graphically, a local complementation on vertex $a$ inverts the sub-graph on the neighbors of $a$. The lower left corner of Fig. 3 demonstrates a local complementation at qubit one, inverting the sub-graph consisting of the qubits two, three and four. A $\sigma_{z}$ measurement on qubit $a$ is equivalent to deleting all edges incident to qubit $a$, and a $\sigma_{y}$ measurement on qubit $a$ corresponds to a local complementation followed by a $\sigma_{z}$ measurement on qubit $a$. Finally, a $\sigma_{x}$ measurement on qubit $a$ is implemented by performing a local complementation on a neighbor $b\in N_{a}$, measuring $a$ in $\sigma_{y}$ basis and performing again a local complementation on $b$. In Fig. 3 possible local Pauli basis measurements on a graph state and their results are summarized.
A.2 State Extraction Methods
It was shown in [49] that it is in general NP-complete to decide if a graph state can be extracted from a given graph state by local Clifford operations, local Pauli measurements and classical communication. Several methods have however been discussed to extract specific target states from resource states in the past. We review two approaches to extract single Bell-states from a given resource state, where we refer to the first one as the isolation strategy and the second one as the X-protocol [31].
A.2.1 Isolation strategy
The idea of the isolation strategy boils down to the isolation of the shortest path between the Bell-state constituents by performing Pauli $Z$ basis measurements on the neighbors. To establish the Bell-state between the end vertices one measures all intermediate qubits along the isolated path using Pauli Y or Pauli X measurements. This strategy has been proposed by Raussendorf et al. [34] to isolate (circuit) wires in a measurement-based quantum computer. This approach comes with several disadvantages. First, it necessitates numerous measurements in the Pauli Z basis to isolate the path, introducing noise to the cluster state. Second, it fails to preserve connectivity (in terms of entanglement) within the cluster, as Pauli Z measurements essentially cut holes into the cluster state. The latter might be problematic, especially if the goal of a protocol corresponds to having multiple, parallel Bell-states.
A.2.2 X-Protocol
An alternative protocol to extract Bell-states from a cluster state has been proposed by Hahn et al. [31], referred to as X-protocol. It first measures all qubits along the shortest path in the Pauli $X$ basis and then removes all qubits still adjacent to the Bell-state by Pauli $Z$ measurements. Ref. [31] shows that the X-protocol requires in the worst case the same number of measurements as the isolation strategy. Note that the butterfly network [50] is the smallest example for the X-protocol. Other works [33, 32] have shown that a staircase-formed measurement path should be preferred over a straight path in a 2D cluster state, because it reduces the remaining Pauli $Z$ basis measurements to the initial neighbors attached to the Bell pair. In the following section we report the potential of the X-protocol for establishing multiplexed communication in a two dimensional cluster state.
Appendix B Proof of Zipper-scheme
In this section we show that the zipper-scheme restores the structure of the remaining cluster state while creating a Bell-state between the two end qubits. We use the graphical and set theoretical rules for Pauli measurements and local Clifford operations from the Refs. [46, 45]. We consider a 2D cluster state where we apply the zipper-scheme along a diagonal path of qubits, labeled from $v_{1}$ to $v_{n}$, to generate a Bell-state between the two end qubits $b_{1}$ and $b_{2}$ of the path, see also Fig. 4. The measurements on the diagonal impact the neighborhood along the path. We distinguish two kinds of qubits in the vicinity of the measured qubits, namely exclusive neighbor qubits that are in the neighborhood of only one measured or the Bell-state qubit, labeled with $e_{i}$ variables, and qubits that are neighbors of at least two measured qubits on the diagonal, labeled with $r_{i}$ variables. The left-most sub-figure in Fig. 4 sketches the situation for a diagonal consisting of six qubits. In the following we explain the zipper-scheme step by step by starting with the Pauli X measurement of $v_{1}$.
Pauli X measurement
Figure 4: Core of zipper-scheme: The aim is to establish the Bell-pair $(b_{1},b_{2})$ with the zipper-scheme on the diagonal $v_{1}$ to $v_{6}$, the initial setting is shown in first step. The second step shows the result of a local complementation on the qubit $b1$, which enlarges the neighborhood of $v_{1}$. The next step shows the Pauli Y measurement on $v_{1}$, which establishes the edge $(b_{1},v_{2})$, necessary for the next measurement, as well as the edge $(r_{1},r_{2})$, necessary for restoring the cluster state, and it removes the qubit $r_{1}$ from the measurement collective, which removes the necessity to measure it in Pauli Z basis. The last step shows the final local complementation on $b_{1}$, which restores a similar configuration as in the initial step, merely $r_{1}$ is replaced by $r_{2}$ and the exclusive neighbors switch from $b_{1}$ to $v_{2}$ and from $v_{1}$ to $b_{1}$.
We recall from section A.1 that a Pauli X measurement on the qubit $v_{1}$ corresponds to a local complementation on a special neighbor (which we choose to be $b_{1}\in N_{v_{1}}$), followed by a Pauli Y measurement on the qubit $v_{1}$ and another local complementation on the special neighbor.
When measuring $v_{1}$ the first step corresponds to a local complementation on the selected, special neighbor, namely on qubit $b_{1}$. This extends the edge set of the sub-graph induced by the neighbors of $b_{1}$ to a fully-connected sub-graph on the vertex set $\{e_{1},e_{2},r_{1},v_{1}\}$, see the second sub-figure in Fig. 4. Note that this also extends the neighborhood of the measurement qubit $v_{1}$, which we virtually measure in the Pauli Y basis next. We note that this virtual measurement corresponds to a local complementation on $v_{1}$ followed by a Pauli Z measurement. The virtual Pauli Y measurement induces many changes to the edge set of the graph, as one can see in the third sub-figure in Fig. 4. For example, the qubit $r_{1}$ which was a neighbor of two qubits along the measurement path gets detached from both qubits $b_{1}$ and $v_{2}$ on the path. This implies that $r_{1}$ is not a direct neighbor neither to the qubits on the diagonal path nor the Bell-state qubits. Furthermore, this circumstance removes the necessity to measure it in the Pauli Z basis. Also, the virtual Pauli Y measurement creates edges between the qubits $r_{1}$ and $r_{2}$ as well as between $b_{1}$ and $v_{2}$, where the latter is the key ingredient to continue with the X-protocol. Finally, note that exclusive neighbors hop from $b_{1}$ to $v_{2}$ and from $v_{1}$ to $b_{1}$. The final step is another local complementation on the special neighbor $b_{1}$, which removes the fully-connected graph of the vertices $\{v_{2},r_{2},e_{3}\}$ from the graph state, see the last sub-figure in Fig. 4. This last step removes the edge between the next qubit to measure $v_{2}$ and $r_{2}$, and thus, it restores a similar setting as for measuring the qubit $v_{1}$. In this setting we find that $r_{2}$ is a neighbor of the Bell-state qubit $b_{1}$ and $v_{3}$ in a similar manner as $r_{1}$ was for $b_{1}$ and $v_{2}$ in the previous step. Specifically, since $r_{2}$ is now in a similar situation as $r_{1}$ was initially the same observations outlined above apply when measuring $v_{2}$. In summary, when we measure the qubits $v_{1}$ to $v_{n}$, $b_{1}$ connects step-by step to the next qubit until it ends up connecting to $b_{2}$. The exclusive neighbors $e_{i}$ toggle between $b_{1}$ and the qubit to be measured $v_{j}$ in each measurement step $j$. Importantly, all qubits $r_{i}$ detach from the measurement path and connect with each other in such a way that they restore the underlying cluster state structure. The exclusive neighbors $e_{i}$ accompany the Bell states and measurement path until the very end which results in the necessity to remove them in the end.
Appendix C Horizontal and vertical transport of multiple data lines in a 2D cluster state
In this section we show how to transport $n$ parallel Bell-states $(A_{1},B_{1}),\dots,(A_{n},B_{n})$ on a straight horizontal or vertical line in a 2D cluster state. To implement such a parallel transport, we first apply the zipper-scheme to diagonally transport the first data line ($A_{1}$) to the last line ($A_{n}$). By using the cluster state restoring property of the zipper-scheme, we transport in a second step the second data line ($A_{2}$) to the second last position ($A_{n-1}$) again via the zipper-scheme. We repeat this until we reversed the order of all lines, namely to $(A_{n},A_{n-1},\dots A_{1})$. Next, we apply the same measurement strategy again. We observe that we recover the original order of the lines through that. However, we note that we must keep one qubit on each permuted line, resulting in $n$ qubits between the two measurement strategies, to have cluster states before starting the second application of the zipper-scheme. The final step is to apply Pauli Y measurements to merge these intermediate qubits on each line in between to the create the final Bell-states. If we have a distance of $N$ qubits between a Bell pair, we can retrieve from a 2D cluster state at most $\mathcal{O}(N/2)$ Bell-states, and note that a fixed number of measurements, scaling with the number of parallel lines, is required to transport parallel lines. In Fig. 5 we show the measurement strategy to transport four Bell-states in parallel step by step. Notice that in contrast to using the isolation strategy, there are no disconnected lines between data lines. This implies (roughly) a factor of two improvement in required resources to transport multiple data lines in parallel.
Pauli X measurement
Pauli Z measurement
repeat measurement pattern on right-hand side
Pauli Y measurement
Figure 5: Measurement pattern for a four qubit parallel transport in horizontal direction. The first three sub-figures show the measurement pattern used to achieve an inversion of order of the input data. The fourth sub-figure repeats the measurement strategy a second time to revert this inversion. The last figure shows the final result after performing the Pauli Y measurements on the intermediate qubits.
Appendix D Taking corners in a 2D cluster state with multiple data lines
In this section we demonstrate how to use the zipper-scheme to change the direction of a quantum data bus from vertical to horizontal (and vice versa). For that purpose, we apply the zipper-scheme (in inward direction) to the outermost data line until we diagonally reach the inner most data line, resulting in a perpendicular turn. We illustrate this step also by the leftmost blue line in Fig. 6. The cluster state preservation property of this first step enables us to apply the zipper-scheme to the second most left qubit until we reach again the innermost qubit. We observe that this results in a perpendicular turn of the second data line of our quantum data bus. Repeatedly applying this pattern results in a full turn of all data lines, as illustrated in Fig. 6.
Pauli X measurement
Pauli Z measurement
Figure 6: The figure shows a step-by-step guide for the measurement pattern to perform a perpendicular turn from horizontal to vertical in a 2D cluster state with four parallel data lines. The zipper-scheme is applied from the left- to right-most qubits, where each application of the zipper-scheme enables the next measurement.
Appendix E Straight to diagonal routing of multiple data lines in a 2D cluster state
In this section we present the measurement pattern to change from parallel to diagonal data lines of a quantum data bus in a 2D cluster state. Applying the zipper-scheme to the last data line in a parallel transport together with the restoration property of the zipper-scheme enables to transport the next qubit along the seam closed by the zipper-scheme. The top part of Fig. 7 depicts the result after switching the direction of the last qubit of a horizontal quantum data bus to diagonal (blue) using the zipper-scheme. We transport the second-last qubit now by measuring along the seam (red line) produced by the zipper-scheme of the last qubit in the first step. The figure at the bottom of Fig. 7 shows the result after the both qubits of the quantum data bus switched the direction from horizontal to diagonal. Note that the length of the initial path for the zipper-scheme determines the number of qubits for the transfer, where a length of $2n$ enables to transfer $n$ qubits. Furthermore, we note that this approach inverts the order of data lines in the quantum data bus, similar as for parallel transport (see also Appendix C for more information).
Pauli X measurement
Pauli Z measurement
Figure 7: Measurement pattern to change between parallel and diagonal information transport is shown. The qubit in the bottom of the parallel line stack has been transported first by with the zipper-scheme. The next qubit in the stack is transported by applying the zipper-scheme along the seam, produced by the first application of the zipper-scheme.
Appendix F ’V’ turn of multiple data lines in a 2D cluster state
In this section we demonstrate the measurement pattern to perform a ’V’ turn of parallel data lines. It basically boils down to applying the same measurement strategy as in the previous section of Appendix E, i.e. to change from straight to diagonal transport, just twice. In the top part of Fig. 8 we applied the zipper-scheme along the blue, dashed paths in order to propagate the bottom data line to the top. We use the closed seam (red) of the zipper-scheme, like for straight-to-diagonal transport, to propagate the next data line to the top. To propagate the qubits now from top to bottom again, we apply the previous measurement pattern for all data lines again, but now downwards in such a way that the resulting paths correspond to a ’V’. We note that we require one intermediate qubit per line at the turning point to enable the full functionality of the zipper-scheme. In the bottom part of Fig. 8 the intermediate result for a ’V’ turn of two parallel data lines to the top and one to the bottom is shown. In order to turn around $n$ data lines, a total number of $2(n\times n)$ is required.
Pauli X measurement
Pauli Z measurement
Figure 8: Measurement strategy to perform a ’V’ turn on two parallel data lines. Top sketch illustrates the application of the zipper-scheme along the blue, dashed path to propagate the bottom data line to the top, and the next data line will be propagated to the top by applying the zipper-scheme along the red seam created the zipper-scheme. The bottom part sketches the ’V’ turn after both data lines have been brought up and the bottom line has been brought down. As a final step the two qubits in the middle have to be merged by Pauli X measurements.
Appendix G Fade-in and -out to a quantum data bus
We achieve a fade-in and -out to or from a quantum data bus by placing Pauli Z measurements at appropriate positions. In Fig. 9 we show a fade-out of a three qubit data line bundle where the yellow measurements correspond to Pauli Z measurements for cutting out the data lines, which are merged subsequently by the orange Pauli X measurements.
Pauli X measurement
Pauli Z measurement
Figure 9: Pauli Z measurements are depicted by the yellow badges, and those cut out the measurement path. By measuring the qubits along the orange lines in the Pauli X basis, we achieve the fade-out of the three data lines.
Appendix H Zipper-scheme for one-dimensional cluster state and GHZ state extraction
In this section, we show how to apply the zipper-scheme to extract one-dimensional cluster states and GHZ states from a two dimensional cluster state. We point out that by using insights from Ref. [51] together with our strategy we generate GHZ states. The simplest way to extract a linear cluster state from a two dimensional cluster state is by applying the zipper-scheme and not measuring qubits on the path. In the first part from the left of Fig. 10 we show how to extract a GHZ state or a three qubit linear cluster state from a 2D cluster state by applying the zipper-scheme along the purple measurement path. However, we note that this comes with an increased number of qubits we have to measure in the Pauli Z basis (yellow qubit along the seam), because (green) edges to the target state remain in addition to the (red) edges due to the zipper-scheme. Overall, the additional Pauli Z measurements of qubits reduce the connectivity in the remaining resource state.
Pauli X measurement
Pauli Z measurement
Figure 10: Linear cluster or GHZ extraction: In the first sub-figure we apply the zipper-scheme along a (purple, dashed) staircase-shaped path to extract a one-dimensional cluster or GHZ state (purple qubits) from the 2D cluster state by keeping intermediate qubits unmeasured, and measuring additional qubits (yellow) within the Pauli Z basis. In the middle we demonstrate that the remaining cluster state, due to the zipper-scheme, enables to extract two one-dimensional cluster or GHZ state adjacent to each other. In the right part an additional one-dimensional cluster or GHZ state (turquoise) is extracted orthogonal to two established states.
H.1 Crossing of 1D cluster states and GHZ states
Choosing the locations of the qubits in a clever way enables us to extract multiple linear cluster states by measuring along a staircase path, similarly as discovered for Bell states. In particular, we split the staircase into two sub-staircase segments such that the beginning and the ending of each stair segment look in the opposite direction. From that condition follows directly, that half of the staircase qubits cannot be chosen, if we want to measure parallel staircases directly adjacent. Furthermore, if we choose the remaining qubits at the same position for all parallel staircase paths, we only have to remove a single exclusive neighbor for each qubit we keep in the linear cluster state. In the middle part of Fig. 10 two parallel linear cluster or GHZ states of size three are extracted by two sub-staircase measurement paths from a two dimensional cluster state, whereas the yellow qubits need to removed from the target states by Pauli Z measurements.
Moreover, we can still extract further linear cluster or GHZ states orthogonal to first ones from the resulting resource state, but the total number of crossing linear cluster states reduces by the number qubits unmeasured on the path. In the right part of Fig. 10 we show a single linear cluster or GHZ state (turquoise) crossing the two already established ones. Another (yellow) qubits needs to be removed, when we extract the orthogonal linear cluster states.
In summary, using the zipper-scheme to distill a linear cluster or GHZ state from the 2D cluster works, but demands a measurement for each additional qubit we add to the Bell pair. Additionally, the generation of GHZ states [51] with more than three qubits from the linear cluster state requires approximately half of the qubits to measured.